# The evolution of cooperation with Q-learning: the impact of information perception

### v1

@@ IV.1 Stage I — Cooperation emergence
Sub-stage ITwo novices both prefer defection resulting in the combination of (D,D).
@@
At the beginning, both players are unfamiliar with the environment, thus they prioritize immediate payoffs and learn that D is more beneficial, leading to $\Delta Q_{D}^{M,Y}<0$. Therefore, both exhibit self-interested behavior in the form of mutual defection (D, D). However, the action preference combinations of (D, D) bring very low payoffs to both parties, which weakens the advantage of choice D and cause both values of $\Delta Q_{D}^{M,Y}$ back to zero [Fig. 5(a)]. The intermittent decreases are due to the action of C by exploration. When one of the values of $\Delta Q_{D}^{M,Y}\rightarrow 0$, meaning that their preferences in D are about to change.
@@
Sub-stage IIPlayer M’s action preference shift leads to a new action preference combination (C, D) $\leftrightarrow$ (D, C).
@@
When $\Delta Q_{D}^{M}>0$ [Fig. 5(a)], the player M’s action preference is shifted from D to C. Correspondingly, the state of the system also undergoes the same change, and the system then enters a new action preference combination (C, D) $\leftrightarrow$ (D, C) [Fig. 5(b)]. However, this action preference combination fails to persist in the presence of exploration.
@@
Sub-stage IIIExploratory behavior of both parties favors cooperation and mutual cooperation (C, C) is formed.
@@
Within the action preference combination (C, D) $\leftrightarrow$ (D, C), the exploratory behavior of both parties is conducive to the growth of the utility function $Q_{{s_{l}},{C}}$ [Fig. 5(b)], and the values of $Q_{C,C}^{M}$ and $Q_{D,C}^{Y}$ increase discontinuously. Correspondingly, $\Delta Q_{C}^{M}$ and $\Delta Q_{D}^{Y}$ show an increasing trend [see inset in Fig. 5(b)], indicating a gradual shift towards cooperation. Due to asymmetric information causing a faster increase in $\Delta Q_{C}^{M}$, see Appendix A, individual M first transitions to cooperation in state D, leading the system to enter state C and establishing a stable positive feedback loop of mutual cooperation. As a result, the system enters a new action preference combination (C,C), the value of $Q_{C,C}^{M,Y}$ remains unchanged after continuous rise in Fig. 5(c).
@@
Sub-stage IVAsymmetric information leads to exploitation of individual M by individual Y, the action preference combination (C,D) is formed.
@@
The action preference combination (C,C) remains unstable. The exploration behavior – defection of both players leads to an increase in $Q_{C,D}^{M,Y}$. However, due to asymmetric information, $Q_{C,D}^{Y}$ increases more rapidly, and $\Delta Q_{C}^{Y}$ is falling at a faster rate than $\Delta Q_{C}^{M}$ [see inset in Fig. 5(c)]. For more details, see Appendix A. Consequently, player Y transitions from cooperation to defection in state C first, leading the system enter the action preference combination (C, D). This combination can be viewed as a process of exploitation and tolerance. For individual M, positive feedback from prior mutual cooperation results in $\Delta Q_{C}^{M}>0$, making M inclined to cooperate even when faced with defection, showing tolerance. Thus, individual Y can exploit M by choosing defection for a period.
@@
Sub-stage VPlayer M implements a punishment-like policy on player Y, the corresponding action preference combination is (D,D) $\leftrightarrow$ (C,D)
@@
However, tolerance within the action preference combination (C, D) is limited. Frequent exploitation by the opponent causes a continuous decline in $Q_{C,C}^{M}$ [Fig. 5(d)] and $\Delta Q_{C}^{M}$ to show a decreasing trend [see inset in Fig. 5(d)]. When $\Delta Q_{C}^{M}<0$, individual M switches from action C to D in state C, transitioning the system to the combination preference of (D,D) $\leftrightarrow$ (C,D). Within this combination, individual Y’s persistent exploitation from the previous sub-stage becomes intermittent, resulting in a reduced payoff and causing $Q_{C,D}^{Y}$ to start declining [Fig. 5(e)]. This can be seen as a punishment process by individual M towards individual Y. This process causes $\Delta Q_{C}^{Y}$ to rise [see inset in Fig. 5(e)]. When $\Delta Q_{C}^{Y}>0$, individual Y reverts to cooperation, forming a positive feedback loop that returns the system to (C,C).
@@
Figure 6: Cooperation collapse in Stage II. (a) The time evolution of $\Delta Q_{C}^{Y,M}$. The upper panel shows the time evolution of $\Delta Q_{C}^{Y}$, where intermittent occurrences of $\Delta Q_{C}^{Y}<0$ can be observed, indicating the accumulation of exploitation of individual M by individual Y. The lower panel shows the evolution of $\Delta Q_{C}^{M}$, with corresponding intermittent oscillations observed in the upper panel, each oscillation representing a punishment process of individual Y by individual M. (b) The time evolution of $Q_{D,C}^{M}$. It can be observed that the decrease in $Q_{D,C}^{M}$ mainly occurs during the punishment process of individual Y by individual M, with the corresponding action preference combination being (D,D)$\leftrightarrow$(C,D). The inset shows the evolution of $Q_{D,C}^{M}$ over different time periods. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ IV.2 Stage II — Cooperation collapse
In Stage I, we observe that individual M frequently “forgives” individual Y and re-established mutual cooperation. However, an intriguing phenomenon emerges afterwards: individual M gradually loses patience and is no longer inclined to cooperate. As shown in Fig. 6(b) and the inset, the decline in $Q_{D,C}^{M}$ primarily occurs during sub-stage V. This indicates that each time individual M punishes individual Y, M’s inclination to choose cooperation in state D diminishes. Once tolerance is completely eroded, the system transitions into state D, resulting in a collapse of cooperation. Consequently, when the opponent exploits again, the system shifts to a state of mutual defection (D,D).
Figure 7: Cooperation reestablishment in Stage III. It shows the evolution of action combination preferences, and the temporal evolution of $Q_{s_{l},a_{m}}^{M,Y}$ or $\Delta Q_{s_{l}}^{M,Y}$. Here, the action preference combination (C,C)(D,D) indicates that individual M chooses defect in state C and cooperate in state D, causing the system to cycle between (C,C)$\leftrightarrow$(D,D). The same applies to (D,D)(C,D). The action pair (D,D)(D,C) indicates that in state D, individual Y alternates between cooperation and defection. Similarly, (D,D)(C,D) shows that individual M switches between cooperation and defection in state C. The sudden declines in (a) are because of occasional cooperation by exploration, where the action of defection brings to a reward $\pi=1+b$. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ IV.3 Stage III — Cooperation reestablishment
There the system transitions away from (D,D) to (C,C) again. Three distinct sub-stages can be divided in this stage.
Sub-stage ISimultaneous cooperative exploration breaks mutual defection, triggers (C,C) $\leftrightarrow$ (D,D) cyclic state.
Within the mutual defection state, the payoff $\pi$ for either is zero, reducing their preference in defection. This is evidenced by the upward trend in $\Delta Q_{D}^{M,Y}$ shown in Fig. 7(a), the intermittent declines are due to occasional cooperative actions during exploration. Unilateral cooperation, however, only strengthens the other player’s preference for defection because their preference in state C remains defection (i.e., $\Delta Q_{C}^{M,Y}<0$). Simultaneous cooperation by both players can alter this situation. When both choose to cooperate, they each receive a payoff $\pi=R$, which triggers an increase in $Q_{D,C}^{M,Y}$ and leads to $\Delta Q_{D}^{M,Y}>0$, indicating a reversal in preference as shown in Fig. 7(b). The system then enters a cyclical state of (C,C) $\leftrightarrow$ (D,D). However, this action preference combination cannot be sustained under weak exploration.
@@
Within the action preference combination (C,C) $\leftrightarrow$ (D,D), the exploration behavior – defection of both players leads to an increase in $Q_{D,D}^{M,Y}$. Due to asymmetric information, $Q_{D,D}^{M}$ increases more rapidly (for more details, see Appendix A. ), causing $\Delta Q_{D}^{M}$ to decrease faster than $\Delta Q_{D}^{Y}$ [see inset in Fig. 7(b)]. Consequently, player M transitions from cooperation to defection in state D first, leading the system enter the action preference combination (D, C) – a process is similar to the exploitation and tolerance observed in sub-stage IV of stage I, with the roles reversed: M exploits Y, while Y tolerates M.
Then, player Y implements a similar punishment-like strategy on player M. Within the action preference combination (D, C), M’s continuous exploitation leads to a persistent decline in $Q_{D,C}^{Y}$. When $\Delta Q_{D}^{Y}\rightarrow 0$, Y gains no advantage in choosing either cooperation or defection, causing $\Delta Q_{D}^{Y}$ to fluctuate around zero [Fig. 7(c)]. The corresponding action preference combination is (D,D) $\leftrightarrow$ (D,C), which then predominantly shifts to mutual defection (D,D) $\leftrightarrow$ (D,D). This process results in an increase in $\Delta Q_{D}^{M}$.
When $\Delta Q_{D}^{M}>0$, individual M re-chooses cooperation in state D. The roles of M and Y then reverse, repeating the previously described process. During this period, fluctuations of $\Delta Q_{D}^{M}$ around zero occasionally revert the system to state C, leading to a decrease in $Q_{C,D}^{M}$ [Fig. 7(d)]. When $\Delta Q_{C}^{M}>0$, the player M’s action preference in state C shifts from defection to cooperation. However, due to the lack of positive returns, $\Delta Q_{C}^{M}$ fluctuates around zero [Fig. 7(e)]. This punishment-like strategy on player Y again results in an increase in $\Delta Q_{C}^{Y}$. When $\Delta Q_{C}^{Y}>0$, Y’s action preference in state C also shifts towards cooperation. This indicates that once the system returns to state C, the positive feedback from mutual cooperation can re-establish cooperation between both parties.
@@
Up to this point, the two individuals have reached a consensus to cooperate in state C. When individual M re-chooses cooperation, the system enters state C and mutual cooperation is successfully re-established. In Fig. 7(f), a result identical to that in Fig. 5(c) indicates that the system returns to the Stage I evolution process.
@@
Finally, to gain an intuitive understanding of the cooperation evolution in Scheme III, we show evolutionary paths for four typical dilemma strengths $b$, see Fig. 8. These shown are obtained by the following procedures. Starting with all possible combinations of the two Q-tables (i.e. $4\times 4$ cases), we monitor the evolution of these combinations, where some “attractors” are observed. For a small value of dilemma strength ($b=0.1$), mutual cooperation is the only stable solution, while for a large value ($b=0.4$) mutual defection is exclusively stable. For the cases in between ($b=0.2,0.3$), the two attractors compete with each other, the evolution of the system is up to which basin of attraction are their initial conditions located. The observations align with the overall picture discussed above.
Figure 9: (Color online) Phase diagram. The color-coded averaged cooperation preference $f_{c}$ in the domain ($\gamma$, $\alpha$), (a-c) are respectively for Scheme I - III. The red regions indicate that cooperation dominates, which often emerge for the combination of a small the learning rate $\alpha$ and a large the discount factor $\gamma$. Each data is averaged 100 realizations, and for each realization the data is averaged 500 rounds after a transient of $2\times 10^{8}$. Other parameters : $\epsilon$ = 0.01, $b$ = 0.2.
V 5. Further comparison
@@
In this section, we first show the phase diagrams for three different schemes in Fig. 9, which reports the average cooperation preference $f_{c}$ in the two key learning parameters domain ($\gamma$, $\alpha$) by fixing the dilemma strength $b=0.2$. We find that in Scheme I, there is no emergence of cooperation across the region for the given $b$ [Fig. 9(a)], instead decent levels of cooperation are observed in the other two schemes. As shown in Fig. 9(b, c), the red regions indicate that cooperation dominates ($f_{c}\sim 0.8$), where the learning rate $\alpha$ is mostly small and the discount factor $\gamma$ is large. The observation can be interpreted as that a high level of cooperation emerges only when players both pay attention to their historical experiences and have a long-term vision. Besides, the region dominated by cooperation within the Scheme III is wider than that of Scheme II. Detailed examination shows that in the case of asymmetric information, a moderate degree of future expectation $\gamma$ is sufficient to trigger the emergence of cooperation given a small value of learning rate $\alpha$.
@@
Apart from the average cooperation preference $\langle f_{c}\rangle$ at the final state, the coverage time towards the final state also matters. Fig. 10(a) shows that average converge time $t_{c}$ for the system towards full cooperation are much shorter in Scheme III than the values within Scheme II. Across the whole range of $b$, the converge time in Scheme II is about one order larger compare to the case of Scheme III. In Fig. 10(b), we can observe that there is a crossover in the average cooperation preference as $b$ is varied. A higher $\langle f_{c}\rangle$ in Scheme II is observed when $b<0.26$, while the opposite observation is made when $b>0.26$. The reasons behind the difference shown in Fig. 10 is closed related to the evolutionary mechanism of Scheme II, which is analyzed in the Appendix B.
VI 6. Discussion

### v2

@@ IV.1 Stage i — Cooperation emergence
Sub-stage iTwo novices both prefer defection resulting in the preference combination of $(D,D)_{D}$.
@@
At the beginning, both players are unfamiliar with the environment, thus they prioritize immediate payoffs and learn that D is more beneficial, leading to $\Delta Q_{D}^{M,Y}<0$. Therefore, both exhibit self-interested behavior in the form of mutual defection $(D,D)_{D}$. However, the action preference combinations of $(D,D)_{D}$ bring very low payoffs to both parties, which weakens the advantage of choice D and cause both values of $\Delta Q_{D}^{M,Y}$ back to zero [Fig. 5(a)]. The intermittent decreases are due to the action of C by exploration. When one of the values of $\Delta Q_{D}^{M,Y}\rightarrow 0$, their preferences in D are about to change.
@@
Sub-stage iiPlayer M’s action preference shift leads to a new action preference combination $(C,D)_{D}$leftrightarrow $(D,C)_{C}$.
@@
When $\Delta Q_{D}^{M}>0$ [Fig. 5(a)], the player M’s action preference is shifted from D to C. Correspondingly, the state of the system also undergoes the same change, and the system then enters a new action preference combination $(C,D)_{D}$leftrightarrow $(D,C)_{C}$ [Fig. 5(b)]. However, this action preference combination fails to persist in the presence of exploration.
@@
Sub-stage iiiExploratory behavior of both parties favors cooperation and mutual cooperation $(C,C)_{C}$ is formed.
@@
Within the action preference combination $(C,D)_{D}$leftrightarrow $(D,C)_{C}$, the exploratory behavior of both parties is conducive to the growth of the utility function $Q_{{s_{l}},{C}}$ [Fig. 5(b)], and the values of $Q_{C,C}^{M}$ and $Q_{D,C}^{Y}$ increase discontinuously. Correspondingly, $\Delta Q_{C}^{M}$ and $\Delta Q_{D}^{Y}$ show an increasing trend [see inset in Fig. 5(b)], indicating a gradual shift towards cooperation. Due to asymmetric information causing a faster increase in $\Delta Q_{C}^{M}$ [see Appendix A], individual M first transitions to cooperation in state D, leading the system to enter state C and establishing a stable positive feedback loop of mutual cooperation. As a result, the system enters a new action preference combination $(C,C)_{C}$, the value of $Q_{C,C}^{M,Y}$ remains unchanged after continuous rise in Fig. 5(c).
@@
Sub-stage ivAsymmetric information leads to exploitation of individual M by individual Y, the action preference combination $(C,D)_{C}$ is formed.
@@
The action preference combination $(C,C)_{C}$ remains unstable. The exploration behavior – defection of both players leads to an increase in $Q_{C,D}^{M,Y}$. However, due to asymmetric information, $Q_{C,D}^{Y}$ increases more rapidly, and $\Delta Q_{C}^{Y}$ is falling at a faster rate than $\Delta Q_{C}^{M}$ [see inset in Fig. 5(c)]. For more details, see Appendix A. Consequently, player Y transitions from cooperation to defection in state C first, leading the system enter the action preference combination $(C,D)_{C}$. This combination can be viewed as a process of exploitation and tolerance. For individual M, positive feedback from prior mutual cooperation results in $\Delta Q_{C}^{M}>0$, making M inclined to cooperate even when faced with defection, showing tolerance. Thus, individual Y can exploit M by choosing defection for a period.
@@
Sub-stage vPlayer M implements a punishment-like policy on player Y, the corresponding action preference combination is $(D,D)_{C}$leftrightarrow $(C,D)_{D}$.
@@
However, tolerance within the action preference combination $(C,D)_{C}$ is limited. Frequent exploitation by the opponent causes a continuous decline in $Q_{C,C}^{M}$ [Fig. 5(d)] and $\Delta Q_{C}^{M}$ to show a decreasing trend [see inset in Fig. 5(d)]. When $\Delta Q_{C}^{M}<0$, individual M switches from action C to D in state C, transitioning the system to the combination preference of $(D,D)_{C}$leftrightarrow $(C,D)_{D}$. Within this combination, individual Y’s persistent exploitation from the previous sub-stage becomes intermittent, resulting in a reduced payoff and causing $Q_{C,D}^{Y}$ to start declining [Fig. 5(e)]. This can be seen as a punishment process by individual M towards individual Y. This process causes $\Delta Q_{C}^{Y}$ to rise [see inset in Fig. 5(e)]. When $\Delta Q_{C}^{Y}>0$, individual Y reverts to cooperation, forming a positive feedback loop that returns the system to $(C,C)_{C}$.
@@
Figure 6: Cooperation collapse in Stage ii. (a) The time evolution of $\Delta Q_{C}^{Y,M}$. The upper panel shows the time evolution of $\Delta Q_{C}^{Y}$, where intermittent occurrences of $\Delta Q_{C}^{Y}<0$ can be observed, indicating the accumulation of exploitation of individual M by individual Y. The lower panel shows the evolution of $\Delta Q_{C}^{M}$, with corresponding intermittent oscillations observed in the upper panel, each oscillation representing a punishment process of individual Y by individual M. (b) The time evolution of $Q_{D,C}^{M}$. It can be observed that the decrease in $Q_{D,C}^{M}$ mainly occurs during the punishment process of individual Y by individual M, with the corresponding action preference combination being $(D,D)_{C}$$\leftrightarrow$$(C,D)_{D}$. The inset shows the evolution of $Q_{D,C}^{M}$ over different time periods. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ IV.2 Stage ii — Cooperation collapse
In Stage i, we observe that individual M frequently “forgives” individual Y and reestablished mutual cooperation. However, an intriguing phenomenon emerges afterwards: individual M gradually loses patience and is no longer inclined to cooperate. As shown in Fig. 6(b) and the inset, the decline in $Q_{D,C}^{M}$ primarily occurs during sub-stage V. This indicates that each time individual M punishes individual Y, M’s inclination to choose cooperation in state D diminishes. Once tolerance is completely eroded, the system transitions into state D, resulting in a collapse of cooperation. Consequently, when the opponent exploits again, the system shifts to a state of mutual defection $(D,D)_{D}$.
Figure 7: Cooperation reestablishment in Stage iii. It shows the evolution of action combination preferences, and the temporal evolution of $Q_{s_{l},a_{m}}^{M,Y}$ or $\Delta Q_{s_{l}}^{M,Y}$. Here, the action preference combination $(C,C)_{D}$ $(D,D)_{C}$ indicates that individual M chooses defect in state C and cooperate in state D, causing the system to cycle between $(C,C)_{D}$$\leftrightarrow$$(D,D)_{C}$. The same applies to $(D,D)_{C}$$(C,D)_{D}$. The action pair $(D,D)_{D}$$(D,C)_{D}$ indicates that in state D, individual Y alternates between cooperation and defection. Similarly, $(D,D)_{C}(C,D)_{C}$ shows that individual M switches between cooperation and defection in state C. The sudden declines in (a) are because of occasional cooperation by exploration, where the action of defection brings to a reward $\pi=1+b$. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ IV.3 Stage iii — Cooperation reestablishment
There the system transitions away from $(D,D)_{D}$ to $(C,C)_{C}$ again. Three distinct sub-stages can be divided in this stage.
Sub-stage iSimultaneous cooperative exploration breaks mutual defection, triggers $(C,C)_{D}$leftrightarrow $(D,D)_{C}$ cyclic state.
Within the mutual defection state, the payoff $\pi$ for either is zero, reducing their preference in defection. This is evidenced by the upward trend in $\Delta Q_{D}^{M,Y}$ shown in Fig. 7(a), the intermittent declines are due to occasional cooperative actions during exploration. Unilateral cooperation, however, only strengthens the other player’s preference for defection because their preference in state C remains defection (i.e., $\Delta Q_{C}^{M,Y}<0$). Simultaneous cooperation by both players can alter this situation. When both choose to cooperate, they each receive a payoff $\pi=R$, which triggers an increase in $Q_{D,C}^{M,Y}$ and leads to $\Delta Q_{D}^{M,Y}>0$, indicating a reversal in preference as shown in Fig. 7(b). The system then enters a cyclical state of $(C,C)_{D}$leftrightarrow $(D,D)_{C}$. However, this action preference combination cannot be sustained under weak exploration.
@@
Within the action preference combination $(C,C)_{D}$leftrightarrow $(D,D)_{C}$, the exploration behavior – defection of both players leads to an increase in $Q_{D,D}^{M,Y}$. Due to asymmetric information, $Q_{D,D}^{M}$ increases more rapidly (for more details, see Appendix  A), causing $\Delta Q_{D}^{M}$ to decrease faster than $\Delta Q_{D}^{Y}$ [see inset in Fig. 7(b)]. Consequently, player M transitions from cooperation to defection in state D first, leading the system enter the action preference combination $(D,C)_{D}$ – a process is similar to the exploitation and tolerance observed in sub-stage iv of stage i, with the roles reversed: M exploits Y, while Y tolerates M.
Then, player Y implements a similar punishment-like strategy on player M. Within the action preference combination $(D,C)_{D}$, M’s continuous exploitation leads to a persistent decline in $Q_{D,C}^{Y}$. When $\Delta Q_{D}^{Y}\rightarrow 0$, Y gains no advantage in choosing either cooperation or defection, causing $\Delta Q_{D}^{Y}$ to fluctuate around zero [Fig. 7(c)]. The corresponding action preference combination is $(D,D)_{D}$leftrightarrow $(D,C)_{D}$, which then predominantly shifts to mutual defection $(D,D)_{D}$leftrightarrow $(D,D)_{D}$. This process results in an increase in $\Delta Q_{D}^{M}$.
When $\Delta Q_{D}^{M}>0$, individual M re-chooses cooperation in state D. The roles of M and Y then reverse, repeating the previously described process. During this period, fluctuations of $\Delta Q_{D}^{M}$ around zero occasionally revert the system to state C, leading to a decrease in $Q_{C,D}^{M}$ [Fig. 7(d)]. When $\Delta Q_{C}^{M}>0$, the player M’s action preference in state C shifts from defection to cooperation. However, due to the lack of positive returns, $\Delta Q_{C}^{M}$ fluctuates around zero [Fig. 7(e)]. This punishment-like strategy on player Y again results in an increase in $\Delta Q_{C}^{Y}$. When $\Delta Q_{C}^{Y}>0$, Y’s action preference in state C also shifts towards cooperation. This indicates that once the system returns to state C, the positive feedback from mutual cooperation can reestablish cooperation between both parties.
Figure 9: (Color online) The color-coded averaged cooperation preference $f_{c}$ in the domain ($\gamma$, $\alpha$). (a-c) are respectively for Scheme I-III. The red regions indicate that cooperation dominates, which often emerge for the combination of a small learning rate $\alpha$ and a large discount factor $\gamma$. Each data is averaged 100 realizations, and for each realization the data is averaged 500 rounds after a transient of $2\times 10^{8}$ steps. Other parameters : $\epsilon$ = 0.01, $b$ = 0.2.
@@
Up to this point, the two individuals have reached a consensus to cooperate in state C. When individual M re-chooses cooperation, the system enters state C and mutual cooperation is successfully reestablished. In Fig. 7(f), a result seemingly identical to that in Fig. 5(c) indicates that the system returns to the Stage i evolution process.
@@
Finally, to gain an intuitive understanding of the cooperation evolution in Scheme III, we show evolutionary paths for four typical dilemma strengths $b$, see Fig. 8. These are obtained by the following procedures. Starting with all possible combinations of the two Q-tables (i.e., $4\times 4$ cases), we monitor the evolution of these combinations, where some “attractors” are observed. For a small value of dilemma strength ($b=0.1$), mutual cooperation is the only stable solution, while for a large value ($b=0.4$) mutual defection is exclusively stable. For the cases in between ($b=0.2,0.3$), the two attractors compete with each other, the evolution of the system is up to which basin of attraction are their initial conditions located. The observations align with the overall picture discussed above.
V Further comparison
@@
In this section, we first present the average cooperation preference $f_{c}$ in the domain of two key learning parameters ($\gamma$, $\alpha$) for three different schemes, with the dilemma strength fixed at $b=0.2$, as shown in Fig. 9. We find that in Scheme I, there is no emergence of cooperation across the region for the given $b$ [Fig. 9(a)], instead decent levels of cooperation are observed in the other two schemes. In Fig. 9(b, c), the red regions indicate that cooperation dominates ($f_{c}\sim 0.8$), where the learning rate $\alpha$ is mostly small and the discount factor $\gamma$ is large. The observation can be interpreted as that a high level of cooperation emerges only when players both pay attention to their historical experiences and have a long-term vision. Besides, the region dominated by cooperation within Scheme III is wider than that of Scheme II. Detailed examination shows that, in the case of asymmetric information, a moderate degree of future expectation $\gamma$ is sufficient to trigger the emergence of cooperation given a small value of the learning rate $\alpha$.
@@
Apart from the average cooperation preference $\langle f_{c}\rangle$ at the final state, the convergence time towards the final state also matters. Fig. 10(a) shows that average convergence time $t_{c}$ for the system towards full cooperation are much shorter in Scheme III than the values within Scheme II. Across the whole range of $b$, the converge time in Scheme II is about one order larger compare to the case of Scheme III. In Fig. 10(b), we can observe that there is a crossover in the average cooperation preference as $b$ is varied. A higher $\langle f_{c}\rangle$ in Scheme II is observed when $b<0.26$, while the opposite observation is made when $b>0.26$. The reason behind the difference shown in Fig. 10 is closed related to the evolutionary mechanism of Scheme II, which is analyzed in the Appendix C.
VI Discussion